# Perspective Transition of Large Language Models for Solving Subjective Tasks

### v1

@@ Abstract
11footnotetext: Equal contribution.11footnotetext: Corresponding authors.
@@ 1 Introduction
Large language models (LLMs) have exhibited substantial advancements Brown et al. (2020); OpenAI (2023, 2022); Touvron et al. (2023); Jiang et al. (2023) in recent years, demonstrating remarkable performance across various tasks such as mathematical reasoning  Luo et al. (2023); Yang et al. (2023), code generation  Chen et al. (2021); Roziere et al. (2023), and commonsense question answering  Talmor et al. (2019). Meanwhile, research in the realm of subjective tasks remains relatively nascent  Rottger et al. (2022); Kanclerz et al. (2023); Sun et al. (2023). Unlike objective tasks, which are typically well-defined and directly solvable, subjective tasks such as metaphor recognition  Mohler et al. (2016) and dark humor detection  Meaney et al. (2021) require an understanding of context, linguistic subtleties, and varying individual perspectives. These elements are challenging to quantify or measure objectively, thus presenting significant obstacles for current LLMs  Jentzsch and Kersting (2023); Wachowiak and Gromann (2023); Mao et al. (2024).
Figure 1: An example for showing challenges of solving subjective tasks using chain-of-thought prompting.
@@
Chain-of-thought (CoT) prompting style methods have been widely used to elicit the reasoning ability of LLMs  Wei et al. (2022); Kojima et al. (2022); Zhang et al. (2023); Yao et al. (2023). However, in subjective-leaning questions, it is difficult to find a chain-of-thought pathway similar to that in conventional reasoning tasks. Also, due to the nature of the subjective tasks, manually handwriting the reasoning paths for subjective tasks is more challenging and less consistent. Therefore, directly using CoT prompting techniques may not be practical for subjective tasks. Furthermore, the generated reasoning pathways can even mislead the model to provide incorrect answers, as example shown in Figure 1. Therefore, we are motivated to propose a general method to enhance the ability of LLMs to solve various subjective tasks.
In this paper, in contrast to the common way of using LLMs is to let it directly answer the questions based on the LLMs’ own direct perspective (e.g., zero-shot  Brown et al. (2020) or zero-shot-CoT reasoning  Kojima et al. (2022), we propose leveraging different perspectives to better address the aforementioned challenging subjective tasks, inspired by the domain Theory of Mind  Premack and Woodruff (1978); Wellman et al. (2001), which refers to the ability to attribute mental states to oneself and others and to understand that these mental states can influence behavior. Our work is also related to the development of LLM-based multi-agents  Xi et al. (2023); Wang et al. (2024a), where we aim to elicit the capacity of LLMs to understand contexts, analysis problems, and give solutions beyond a single fixed response based on direct perspective.
@@
Given the breadth and complexity of subjective tasks, we propose a Reasoning through Perspective Transition (RPT) method to dynamically select suitable perspectives to solve specific problems. In particular, we consider categorizing current reasoning methods of LLMs into three perspectives, including: 1) direct perspective, which involves the model directly answering questions or tasks based on its internal understanding without considering external factors or alternative viewpoints, 2) role perspective, which focuses on assigning specific roles to the model, simulating different viewpoints or expertise within a given context or scenario, and 3) third-person perspective, which involves the model considering external viewpoints or perspectives beyond its own, similar to how a third party or observer might view a situation. The examples of three different perspectives when solving problems are shown in Figure 2.
@@
To facilitate the dynamic perspective transition during reasoning, we follow the in-context learning  Brown et al. (2020) approach to provide templates for answering from different perspectives through demonstrations, then let the model provide confidence levels  Li et al. (2024); de Vries and Thierens (2024); Bank et al. (2019) for answers to specific questions, and finally answer based on the perspective with the highest confidence. In this manner, our method can freely select among three different perspectives to handle various subjective tasks. This flexibility allows the model to adapt its responses more effectively to nuanced tasks that traditional static methods struggle with. Additionally, this approach is grounded in the hypothesis that LLMs perform better when their operational parameters align with their confidence levels in specific contexts.
@@
To validate the effectiveness of our proposed method, we conduct experiments on four LLMs (including two closed-source models GPT-4  OpenAI (2023)/GPT-3.5  OpenAI (2022), and two open-source models Llama-3  Dubey et al. (2024)/Qwen-2  Yang et al. (2024)) across 12 subjective tasks. Extensive experimental results demonstrates that, compared to previous methods based on a single perspective or some simple ensemble-based methods, our approach can improve the performance consistently.
@@ 2 Related Work
In-Context Learning of LLMs. As the number of model parameters increases, the in-context learning  Brown et al. (2020) ability of LLMs becomes stronger, significantly enhancing the zero-shot and few-shot reasoning capabilities without model fine-tuning. In particular, methods based on chain-of-thought  Wei et al. (2022); Kojima et al. (2022) prompting are widely used. These works aims to elicit the reasoning capability of LLMs through adding the reasoning pathways. However, recent research has shown that such reasoning pathways are mainly effective for math and symbolic reasoning  Sprague et al. (2024). Our work also relies on in-context learning, however, we propose a method based on dynamic perspective transition to elicit knowledge from the different perspective of LLMs, which does not rely on a single reasoning pathway and achieve better results on a wider range of subjective tasks.
Perspective Transition of LLMs. There are various ways to use LLMs currently that are based on different perspectives: 1) Direct prompting methods  Brown et al. (2020); Wei et al. (2022); Kojima et al. (2022) let the model to provide answers based on the factual knowledge or reasoning ability by LLMs themselves directly, without setting specific roles. 2) By assigning roles  Xu et al. (2023); Wang et al. (2024d); Wilf et al. (2024) such as experts and engaging in role-playing dialogue, the internal knowledge of LLMs on specific roles can be elicited. 3) By constructing scenarios through multi-agent cooperation  Wang et al. (2024e, b), debates  Du et al. (2024), or dialogues  Wang et al. (2024c), and then providing answers from a third-person perspective by incorporating contextualized information by the constructed agents. The previous methods only consider a fixed perspective and validate the effectiveness in certain problems. In contrast, through our proposed RPT method based on in-context learning, LLMs are able to adaptively select the most suitable perspective to solve various subjective tasks, which has not been studied in previous research.
@@ 3 Method
The overall pipeline of the proposed RPT is structured into three steps. Firstly, we input the task description and a specific question, prompting the model to select the most appropriate perspective for answering the question. Secondly, the model evaluates and ranks these perspectives based on their confidence levels in addressing the question. Thirdly, the model adopts the perspective with the highest confidence to formulate and deliver the definitive answer.
@@
Step 1: Exploring Diverse Available Perspectives. We first let LLM $\mathcal{M}$ explore diverse perspectives $\mathcal{P}$ according to the description $\mathcal{D}$ and question $\mathcal{Q}$. Specifically, we have:
@@
Step 2: Ranking Perspectives by Confidence Level. Then, base on the perspectives, we let LLM $\mathcal{M}$ list all the perspectives by the confidence level $\mathcal{C}$ in solving the question:
@@ 4.1 Settings
Table 3: Main results of baselines and our proposed RPT method in zero-shot settings. Random represents the result of random prediction with uniform probability, and Majority represents the result of predicting the label with the highest proportion. S1: single direct perspective, S2: single role perspective, S3: single third-person perspective, E: ensemble-based method. D: dynamic perspective. For each dataset, the best result is in bold and the second-best result is underlined.
@@
Single Third-Person Perspective. SPP  Wang et al. (2024e) (solo performance prompting) proposes solo performance prompting by involving multi-turn collaboration with multi-persona. RiC  Wang et al. (2024c) (reason in conversation) first lets model generating dialogues between simulated roles, and then summarize conversations and give final answers according to the additional information from conversations.
Ensemble-based Methods. Ensemble  Messuti et al. (2024); Agrawal et al. (2024) involves combining multiple model generation to enhance prediction accuracy and robustness. Reranking  Farinhas et al. (2024); Kim et al. (2024) reorder different generation options based on requirements and select the optimal result. CoT-SC  Wang et al. (2023) enhances performance by sampling diverse chain-of-thought reasoning paths and selecting the most self-consistent answer.
Models. We evaluate our method on both closed-source models including GPT-4  OpenAI (2023) and GPT-3.5  OpenAI (2022), and open-source Llama-3  Dubey et al. (2024) and Qwen-2  Yang et al. (2024) models. In particular, we use the released API versions of gpt-4-0613 and gpt-3.5-turbo-1106 by OpenAI, and open-source Llama-3-8b-instruct and qwen-2-7b-instruct models released in Huggingface hub. We set the decoding temperature as 00 to maintain the reproducibility of the responses generated by LLMs.
@@ 4.2 Zero-shot Results
Subjective reasoning is a challenging task, and the RPT method effectively elicits the capabilities of LLMs in such tasks. Closed-source models like GPT-3.5 and GPT-4 outperform open-source models like Llama-3 and Qwen-2 in subjective tasks, but the performance gap narrows when using the RPT method. This suggests that these models still possess the knowledge required for subjective reasoning, but it has not been effectively elicited during training. By introducing confidence evaluation during LLM reasoning, RPT selects the method best suited to the model’s strengths, effectively eliciting LLM capabilities in subjective tasks, and compensating for the lack of subjective task data during LLM training.
@@ 4.3 Few-shot Results
In Table 4.1, we present the main results in few-shot settings. Similar to the zero-shot results, RPT method achieves the best average performance across different models. For instance, on Llama-3, RPT surpasses CoT-SC, the best-performing baseline, by 2.06 points.
@@ 5.1 Ablation Study
5.2 Analysis on Inference Cost
The inference cost of modern LLMs is crucial. Using the GPT-3.5 model as an example, we represent the inference cost by the length of the response during the reasoning process before producing the final answer. In Figure 7, using the 3-shot GPT-4 experiment as an example, we plot the length-performance relationship for RPT and the baselines. It can be observed that compared to most baselines, RPT achieves the best performance with a smaller inference cost, demonstrating the efficiency of the dynamic perspective selection approach. As shown in Figure 7, we compare the inference cost of the RPT method with its variants that only allow the use of one or two perspectives. The results indicate that the RPT method incurs a cost similar to that of the variants allowing only one or two perspectives in most datasets, without increasing the inference length, and thus the additional cost is minimal. This is because RPT performs method selection and perspective switching, avoiding the substantial inference cost caused by ensembling multiple perspectives.
@@ 5.3 Analysis on the Number of Shots
On one hand, LLMs exhibit greater flexibility when autonomously evaluating confidence and planning the number of shots during reasoning, allowing them to adapt to unique subjective tasks. On the other hand, providing too many examples may lead to increasing the inference cost, raising the risk of over-fitting, and challenging the instruction-following ability of LLMs. Overall, under the majority of settings for each dataset, RPT achieves the best performance, demonstrating its generalization ability and versatility (See Appendix B.3 for full results).
@@ 5.5 Analysis on Keyword Statistics
Figure 9: Keyword statistics of different perspectives in RPT pipeline.
@@ 6 Conclusion
In this paper, we introduce RPT, a novel method that achieves multi-perspective reasoning and integration by exploring diverse perspectives and ranking them based on confidence. Comprehensive experiments conducted on GPT-4, GPT-3.5, Llama-3, and Qwen-2 demonstrate that RPT effectively integrates various perspectives, enhancing the subjective task-solving capabilities of LLMs without significantly increasing inference costs. This work highlights how LLMs can better handle the fluidity of subjective reasoning, even in the absence of nuanced understanding of perspectives or personal biases. Future research directions include integrating additional reasoning perspectives, developing finer-grained and adaptive perspective taxonomies, and extending our method to broader applications.
@@ B.2 More Cases of RPT
B.3 Full Results of Analysis on the Number of Shots
The full results of analysis on the number of shots is shown in Figure 16. The results are consistent across the four datasets.

### v2

@@ Abstract
11footnotetext: Equal contribution. Work done at Tsinghua University. 11footnotetext: Corresponding authors.
@@ 1 Introduction
Figure 1: Two examples for showing challenges of solving subjective tasks using chain-of-thought prompting. Refer to Appendix B.6 for more examples.
@@
Large language models (LLMs) have exhibited substantial advancements Brown et al. (2020); OpenAI (2023, 2022); Touvron et al. (2023); Jiang et al. (2023) in recent years, demonstrating remarkable performance across various tasks such as mathematical reasoning  Luo et al. (2023); Yang et al. (2023), code generation  Chen et al. (2021); Roziere et al. (2023), and commonsense question answering  Talmor et al. (2019). Meanwhile, research in the realm of subjective tasks remains relatively nascent  Rottger et al. (2022); Kanclerz et al. (2023); Sun et al. (2023). Unlike objective tasks, which are typically well-defined and directly solvable, subjective tasks such as metaphor recognition  Mohler et al. (2016) and dark humor detection  Meaney et al. (2021) require an understanding of context, linguistic subtleties, and varying individual perspectives. Although advanced chain-of-thought (CoT) style methods have been widely used to elicit the reasoning ability of LLMs  Wei et al. (2022); Kojima et al. (2022); Zhang et al. (2023); Yao et al. (2023), these approaches primarily focus on “how to think deeper in its own perspective” OpenAI (2024); DeepSeek-AI et al. (2025), while overlooking “from which perspective to think”.
@@
This leads to modern LLMs performing poorly on subjective tasks challenging to be quantified or measured objectively Jentzsch and Kersting (2023); Wachowiak and Gromann (2023); Mao et al. (2024). Moreover, due to the nature of the subjective tasks, identifying a chain-of-thought pathway similar to that in conventional reasoning tasks is difficult Wang et al. (2024c). For instance, the generated reasoning pathways can even mislead the model to provide incorrect answers, as examples shown in Figure 1. Therefore, directly exploring CoT prompting techniques may not be practical for subjective tasks, motivating us to propose a general method to enhance the ability of LLMs to solve various subjective tasks.
@@
Inspired by the theory of mind  Premack and Woodruff (1978); Wellman et al. (2001), which refers to the ability to attribute mental states to oneself and others and to understand that these mental states can influence behavior, we leverage different perspectives to better address the aforementioned challenging subjective tasks rather than directly answering the questions based on the LLMs’ own direct perspective (e.g., zero-shot  Brown et al. (2020) or zero-shot-CoT reasoning  Kojima et al. (2022). As shown in Figure 2, we categorize reasoning methods of LLMs into three perspectives: 1) direct perspective, which involves the model directly answering questions or tasks based on its internal understanding without considering external factors or alternative viewpoints; 2) role perspective, which focuses on assigning specific roles to the model, simulating different viewpoints or expertise within a given context or scenario; 3) third-person perspective, which involves the model considering external viewpoints or perspectives beyond its own, similar to how a third party or observer might view a situation.
@@
Based on the intuition that LLMs perform optimally when their operational parameters align with confidence levels in specific contexts, we introduce Reasoning through Perspective Transition (RPT) method to dynamically select suitable perspectives to solve specific problems. To enable dynamic perspective transitions in reasoning, we adopt in-context learning  Brown et al. (2020), providing templates for multi-perspective answers. The model then evaluates confidence levels  Li et al. (2024); de Vries and Thierens (2024); Bank et al. (2019) and selects the perspective with the highest confidence. This adaptability allows it to handle diverse subjective tasks more effectively than static methods.
@@
We conduct experiments on four LLMs (including two closed-source models GPT-3.5  OpenAI (2022)/GPT-4  OpenAI (2023), and two open-source models Llama-3  Dubey et al. (2024)/Qwen-2  Yang et al. (2024) across 12 subjective tasks. Extensive experimental results demonstrates that, compared to previous methods based on a single perspective or some simple ensemble-based methods, our approach can improve the performance consistently.
@@ 2 Related Work
In-Context Learning of LLMs. As model parameters increase, LLMs gain stronger in-context learning Brown et al. (2020), enhancing zero-shot and few-shot reasoning without fine-tuning. Chain-of-thought (CoT) prompting Wei et al. (2022); Kojima et al. (2022) is widely used to elicit reasoning by adding explicit reasoning steps, emphasizing “how to think deeperOpenAI (2024); DeepSeek-AI et al. (2025) while neglecting “who should think.” However, recent research Sprague et al. (2024) suggests that such reasoning pathways are primarily effective for mathematical and symbolic reasoning. Our work also relies on in-context learning, however, we propose a method based on dynamic perspective transition to elicit knowledge from the different perspective of LLMs, which does not rely on a single reasoning pathway and achieve better results on a wider range of subjective tasks. Our work also relates to LLM-based multi-agents  Xi et al. (2023); Wang et al. (2024a), aiming to enhance LLMs’ ability to understand context, analyze problems, and generate solutions beyond a single fixed perspective.
Perspective Transition of LLMs. There are various ways to use LLMs currently that are based on different perspectives: 1) Direct prompting methods  Brown et al. (2020); Wei et al. (2022); Kojima et al. (2022) let the model to provide answers based on the factual knowledge or reasoning ability by LLMs themselves directly, without setting specific roles. 2) By assigning roles  Xu et al. (2023); Wang et al. (2024e); Wilf et al. (2024) such as experts and engaging in role-playing dialogue, the internal knowledge of LLMs on specific roles can be elicited. 3) By constructing scenarios through multi-agent cooperation  Wang et al. (2024f, b), debates  Du et al. (2024), or dialogues  Wang et al. (2024c), and then providing answers from a third-person perspective by incorporating contextualized information by the constructed agents. The previous methods only consider a fixed perspective and validate the effectiveness in certain problems. In contrast, RPT is orthogonal to improvements such as deep long-chain CoT DeepSeek-AI et al. (2025), enabling LLMs to adaptively select the most suitable perspective to solve various subjective tasks, which has not been studied in previous research.
@@ 3 Method
The overall pipeline of the proposed RPT is structured into three steps, as shown in Figure 3. Firstly, we input the task description and a specific question, prompting the model to select the most appropriate perspective for answering the question. Secondly, the model evaluates and ranks these perspectives based on their confidence levels in addressing the question. Thirdly, the model adopts the perspective with the highest confidence to formulate and deliver the definitive answer.
Step 1: Exploring Diverse Available Perspectives. We first let LLM $\mathcal{M}$ explore diverse perspectives $\mathcal{P}$ according to the description $\mathcal{D}$ and question $\mathcal{Q}$, promoting more comprehensive reasoning. Specifically, we define:
@@
Step 2: Ranking Perspectives by Confidence Level. Then, based on the generated perspectives, we allow the LLM $\mathcal{M}$ to rank them according to their confidence levels $\mathcal{C}$ in solving the question.
@@ 4.1 Settings
Table 3: Zero-shot results. To demonstrate generalizability, we repeat each set of experiments separated by dashed lines on four LLMs independently. Random represents the result of random prediction with uniform probability, and Majority represents the result of predicting the label with the highest proportion. S1: single direct perspective, S2: single role perspective, S3: single third-person perspective, E: ensemble-based method. D: dynamic perspective. For each dataset, the best result is in bold and the second-best result is underlined.
@@
Single Third-Person Perspective. SPP  Wang et al. (2024f) (solo performance prompting) proposes solo performance prompting by involving multi-turn collaboration with multi-persona. RiC  Wang et al. (2024c) (reason in conversation) first lets model generating dialogues between simulated roles, and then summarize conversations and give final answers according to the additional information from conversations.
Ensemble-based Methods. Ensemble  Messuti et al. (2024); Agrawal et al. (2024) involves combining multiple model generation to enhance prediction accuracy and robustness (refer to Appendix B.3). Reranking  Farinhas et al. (2024); Kim et al. (2024) reorder different generation options based on requirements and select the optimal result. CoT-SC  Wang et al. (2023) improves performance by generating diverse chain-of-thought reasoning paths and selecting the most self-consistent answer, thereby enhancing the robustness and reliability.
Models. We evaluate our method on both closed-source models including GPT-4  OpenAI (2023) and GPT-3.5  OpenAI (2022), and open-source Llama-3  Dubey et al. (2024) and Qwen-2  Yang et al. (2024) models. In particular, we use the released API versions of gpt-4-0613 and gpt-3.5-turbo-1106 by OpenAI, and open-source Llama-3-8b-instruct and qwen-2-7b-instruct models released in Huggingface hub. We set the decoding temperature to 00 to ensure deterministic outputs and maintain the reproducibility of the responses generated by LLMs.
@@ 4.2 Zero-shot Results
Subjective reasoning is highly challenging and performs better on powerful LLMs capable of deep reasoning. RPT almost always achieves the best performance on powerful models capable of deep reasoning, such as GPT-3.5 and GPT-4, while being more unstable on weaker open-source models. This suggests that rather than merely enhancing the ability to think from its own perspective, considering different perspectives can further improve LLM performance, especially for well-trained LLMs capable of long-chain deep reasoning. The performance gap between close-source models and open-source models narrows when using the RPT method, given that LLMs still possess the knowledge required for subjective reasoning and not been effectively elicited during training. RPT selects the method best suited to the model’s strengths, effectively eliciting LLM capabilities in subjective tasks, and compensating for the lack of subjective task data during LLM training.
@@ 4.3 Few-shot Results
In Table 4.1, we present the 3-shot results (refer to Section 5.3 for other settings). Similar to the zero-shot results, RPT method achieves the best average performance across different models. For instance, on Llama-3, RPT surpasses CoT-SC, the best-performing baseline, by 2.06 points.
@@ 5.1 Ablation Study
5.2 Analysis on Performance-Inference Cost
Unlike methods such as ensemble and oversampling, which multiply inference costs, RPT aims to improve performance while reducing inference costs.
To estimate the inference cost, we follow Wang et al. (2024c) by representing the inference cost by the length of the response during the reasoning process before producing the final answer. In Figure 7, using the 3-shot GPT-4 as an example as detailed in Appendix B.4, we plot the length-performance relationship for RPT and the baselines. It can be observed that compared to most baselines, RPT achieves the best performance with a smaller inference cost, demonstrating the efficiency of the dynamic perspective selection approach. As shown in Figure 7, we compare the inference cost of the RPT method with its variants that only allow the use of one or two perspectives. The results indicate that the RPT method incurs a cost similar to that of the variants allowing only one or two perspectives in most datasets, without increasing the inference length, and thus the additional cost is minimal. This is because RPT performs method selection and perspective switching, avoiding the substantial inference cost caused by ensembling multiple perspectives.
@@ 5.3 Analysis on the Number of Shots
On one hand, LLMs exhibit greater flexibility when autonomously evaluating confidence and planning the number of shots during reasoning, allowing them to adapt to unique subjective tasks. On the other hand, providing too many examples may lead to increasing the inference cost, raising the risk of over-fitting, and challenging the instruction-following ability of LLMs. Overall, under the majority of settings for each dataset, RPT achieves the best performance, demonstrating its generalization ability and versatility (See Appendix B.8 for full results and analysis).
@@ 5.5 Analysis on Keyword Statistics
Figure 9: Keyword statistics of different perspectives in the RPT pipeline.
@@ 6 Conclusion
In this paper, we introduce RPT, a novel method that achieves multi-perspective reasoning and integration by exploring diverse perspectives and ranking them based on confidence. Comprehensive experiments conducted on GPT-4, GPT-3.5, Llama-3, and Qwen-2 demonstrate that RPT effectively integrates various perspectives, enhancing the subjective task-solving capabilities of LLMs without significantly increasing inference costs. This work highlights how LLMs can better handle the fluidity of subjective reasoning, even in the absence of nuanced understanding of perspectives or personal biases. Future research directions include incorporating more reasoning perspectives, refining adaptive perspective taxonomies, and expanding applications.
@@ B.2 More Cases of RPT
B.3 Comparison with Ensemble-based Methods
Instead of oversampling responses and simply making a selection using model confidence, the core idea of RPT is that LLMs can leverage their internal priors and the given prompt to directly select the perspective with the highest posterior accuracy. Unlike conventional ensembling approaches, RPT assigns a single optimal perspective to each query without attempting inferencing on multiple perspectives per instance before making a decision. This means that:
  • RPT does not ensemble multiple responses.
  • RPT does not significantly increase the generation window like ensembling does.
  • RPT does not rely on oversampling responses before selection.
Instead, RPT possesses the following features:
  • RPT selects a single perspective after a brief internal deliberation based on the input prompt and then conducts inference solely from that perspective.
  • The computational cost of RPT is comparable to using a single perspective.
  • Despite its simplicity, RPT achieves performance comparable to or better than ensemble methods.
For example, given $k$ candidate perspectives, an ensemble method incurs a response length/inference cost of $k\times$ cost, with a performance of p. In contrast, RPT has a response length/inference cost of approximately 1×1$\times$1 × cost, while achieving performance $\geq p$. We believe that LLMs’ internal priors can be leveraged to preemptively plan and evaluate reasoning chains across different perspectives, allowing them to anticipate ensemble outcomes before actually performing ensembling, thus eliminating unnecessary computational overhead. This potential has not been explored before. Analogous to Mixture of Experts (MoE), RPT can be regarded as a form of “Mixture of Perspectives”, where exactly one perspective is activated per inference. Thus, we believe that RPT is a highly novel approach.
Therefore, instead of simply selecting perspective, we propose the dynamic process of shifting between different perspectives when handling various tasks, rather than merely selecting a single perspective. It highlights the following two aspects:
  • Dynamism: The model autonomously transitions from one perspective to another, rather than making a static selection. This process is an inherent decision made by the model itself.
  • Variability: In practice, this transition occurs through mechanisms such as Chain-of-Thought (CoT), role-playing, multi-agent discussions, and other discourse-based approaches, facilitating perspective shifts. Compared to perspective selection, we emphasize the dynamic and evolving nature of this process.
B.4 Details on Shot Numbers
In Section 5.3, we provide experimental results across multiple shot settings, demonstrating that the 3-shot setting yields the best performance. Moreover, the datasets used in this study primarily consist of binary or ternary classification tasks. Conventionally, 3-shot is a common setting for practical inference. Additionally, both the baseline and our proposed method achieve the best overall performance under this setting, making it a reasonable choice for our primary study on inference cost. Therefore, we choose 3-shot setting to conduct analysis on performance-inference cost experiments.
B.5 Estimation of Inference Cost
The purpose of Section 5.2 is to examine the trade-off between inference cost and performance when using different methods with LLMs. Counting the number of output tokens is the most common approach for estimating inference time, as LLMs process input tokens in parallel with relatively low overhead, whereas output generation is auto-regressive and thus dominates inference latency and cost. For common LLMs, the cost of input tokens is lower than output tokens. Including both input and output tokens in the performance-cost scatter plot would introduce bias and deviate from standard practices in the LLM community.
Therefore, we chose to plot a two-dimensional scatter plot of output token count versus performance. Introducing a third variable would unnecessarily complicate the experiment and make it more difficult to draw clear conclusions.
B.6 More Examples on Subjective Tasks
Perspective in language is essentially a matter of deixis, how entities are positioned relative to the speaker (user), the addressee (LLM), and a potential third party.
  • First-person refers directly to the speaker’s own viewpoint, emphasizing subjectivity and experiential immediacy. It forms the core of self-reference and establishes speaker agency. This is the most common LLM reasoning mode, as it activates the model’s cognitive framing, directly connecting it to the question and reasoning process.
  • Second-person focuses on the listener in an interaction, creating direct engagement. It reduces the narrative distance through dialogic framing, commonly used in imperative or persuasive contexts. LLMs can leverage second-person narration to enhance reasoning in ambiguous scenarios by simulating conversational exchange.
  • Third-person provides an external, often omniscient view of the situation or entities involved. In LLM reasoning, this resembles a form of distancing effect, offering broader perspective through heterodiegetic narration.
As shown in the experiment results, our experiments reveal that for subjective tasks, Zero-shot-CoT does not exhibit a significant advantage over Direct Prompting on GPT-3.5 and GPT-4-0613, and even performs slightly worse on Qwen2-7B-Instruct (Direct Prompt: 62.70, Zero-shot-CoT: 61.42), as detailed in Table 4.1. This suggests that unlike objective tasks, subjective tasks may not allow for a universally applicable reasoning pathway. This observation aligns with findings by Wang et al. (2024d), which systematically studies the performance of LLMs on subjective tasks.
Example 1:
  • Question: Identify whether the sentence “her smile was a blinding sun” is a paraphrase of the metaphoric sentence “her smile was wonderful and fulfilling to see”.
  • Chain-of-Thought Reasoning: While both sentences express admiration for her smile, the second sentence lacks the vivid imagery and specific emotional intensity that the first sentence provides. Thus, the second sentence is not an accurate paraphrase of the first. The answer is False.
Example 2:
  • Question: Identify whether the following examples are intended to be a joke (with dark humor) or not “I have a fish that can breakdance! Only for 20 seconds though, and only once”.
  • Chain-of-Thought Reasoning: The phrase is light-hearted and whimsical about a fish that can breakdance, suggesting a humorous intent rather than a dark theme. Therefore, it does not seem to fit the category of dark humor. The answer is Not.
Example 3:
  • Question: The sentence “Joseph is very kind” is a paraphrase of the metaphoric sentence “Joseph has the heart of a lion”.
  • RiC Reasoning: The second sentence suggests that Joseph possesses qualities associated with a lion, while the first sentence only mentions kindness. The answer is False.
B.7 Discussion on Overconfidence and Underconfidence
Regarding the inherent overconfidence and underconfidence phenomena in LLMs, previous research Yin et al. (2023) and our experiments (as shown in Figure 4) indicate that the higher the confidence of an LLM’s response, the more likely it is to be correct. We have further investigated the impact of different confidence levels on data selection (as shown in Table 6) by explicitly providing confidence scores for different perspectives and selecting perspectives based on these confidence levels.
Our experiments also show that automatically selecting the perspective with the highest confidence achieves an average performance of 77.94, significantly outperforming other baseline methods. In contrast, selecting the second-highest confidence perspective results in an average performance of 68.98, which is similar to random selection (69.05). The lowest-confidence selection, however, performs notably worse, with an average performance of only 60.93, indicating that choosing perspectives with the lowest confidence leads to poor outcomes.
In Figure 12, by comparing human evaluation scores with LLM-generated confidence scores after normalization, we find that LLMs’ confidence predictions are generally accurate. As shown in Figure 10, a per-instance comparison demonstrates a high correlation between LLM-predicted confidence and actual accuracy. In future research, providing additional supervision signals to models based on dataset characteristics and input prompt designs may further mitigate overconfidence and underconfidence issues.
B.8 Full Results of Analysis on the Number of Shots
The full results of analysis on the number of shots is shown in Figure 16. The results are consistent across the four datasets. It can also be observed that compared to zero-shot settings, few-shot settings are not always helpful, which is consistent with the findings by DeepSeek-AI et al. (2025). The following factors may contribute to the observed trends:
  • As shown in Figure 7, increasing the number of shots does not always enhance LLM performance. While providing a few examples can help the model handle unfamiliar tasks and follow instructions, excessive shots can lead to diminishing returns or performance degradation. We chose the 3-shot setting because most datasets consist of binary or ternary classification tasks.
  • The length of the input prompt can impact LLM inference. Excessive in-context examples may distract the model, reducing its ability to follow instructions precisely, which can negatively impact performance.
  • The added examples may introduce noise, affecting reasoning performance. For instance, gpt-3.5-turbo-1106 tends to overfit the provided cases, often favoring the answer from the last example presented in the prompt.
  • The inherent openness and uncertainty of subjective tasks may contribute to this phenomenon. Subjective task responses are often ambiguous, and in few-shot settings, methods like Ensemble, Reranking, and CoT-SC may be influenced by the provided examples, potentially introducing biases and limiting generalization. In contrast, zero-shot performance relies on pre-trained knowledge, making it more stable. Ensemble methods are susceptible to errors from individual models, reranking depends on the quality of initial generations, and CoT-SC may reduce reasoning diversity, diminishing its advantage in maintaining consistency.
B.9 Using Small Models as Classifiers
We experiment with using small discriminative models, such as bert-base-uncased, as classifiers. However, due to the complexity of the datasets involved in RPT and the lack of distinctive content features, the classification performance was generally poor. Specifically, we use a large model to run each of the different perspective methods on the training set. For samples where the answer was correct, we labeled them according to a priority order: 1 (Direct), 2 (Role-Play), 3 (Third-Person), and then split the data into training and test sets with an 8:2 ratio. The results are shown in Table 7, with the highest accuracy reaching only 48%.
Hyperparameters Accuracy (%) F1-Score (%)
epoch: 3, lr: 2e-5 36 23
epoch: 4, lr: 2e-5 33 32
epoch: 4, lr: 1e-5 30 28
epoch: 4, lr: 2e-4 47 35
epoch: 4, lr: 5e-4 48 31
epoch: 8, lr: 2e-5 46 35
Table 7: Performance of small model classifier with different training hyperparameters.
These results suggest that using small models is not a feasible approach. Moreover, introducing additional parameters for perspective ranking increases inference overhead. In contrast, RPT leverages the LLM’s own capabilities to assess the suitability of different perspectives on a given input without requiring additional training or parameters, achieving performance gains more efficiently.